Break-even Financing. Lakeland, Inc., is a U.S.‑based MNC with a subsidiary in Mexico. Its Mexican subsidiary needs a one‑year loan of 10 million pesos for operating expenses. Since the Mexican interest rate is 70 percent, Lakeland is considering borrowing dollars, which it would convert to pesos to cover the operating expenses. By how much would the dollar have to appreciate against the peso to cause such a strategy to backfire? (The one‑year U.S. interest rate is 9%.)
Answer #1

By how much would the dollar have to appreciate against the peso to cause such a strategy to backfire?
Answer is 55.9633%

BORROWING IN USD
Total amount borrowed=10 million peso
Let current exchange rate be x pesos per dollar
So dollars borrowed=10/x million USD
Now lets say dollar appreciates by k%
Total amount to be returned=10/x*1.09 million USD or 10/x*1.09*x*(1+k%) million peso

BORROWING IN PESO
Total amount borrowed=10 million peso
Total amount to be returned=10*(1+70%) million peso

FOR THE STRATEGY TO BACKFIRE
10/x*1.09*x*(1+k%)>=10*(1+70%)
=>k>=1.7/1.09-1
=>K>=55.9633%